The systolic blood pressure (BP) of adults in the USA is nearly normally distributed with a mean of 120 and standard deviation of 20 .
Someone qualifies as having Stage 2 high blood pressure if their systolic blood pressure is 160 or higher.
(a) Around what percentage of adults in the USA have stage 2 high blood pressure? Give your answer rounded to two decimal places.
____________ %
(b) Stage 1 high BP is specified as systolic BP between 140 and 160. What percentage of adults in the US qualify for stage 1? Give your answer rounded to two decimal places.
_____________ %
(c) Your doctor tells you that you are in the 30th percentile (e.g. bottom 30%) for blood pressure among US adults. What is your systolic BP? Give your answer
rounded to two decimal places.
______________

Answer by Boreal(15235)

a. This is asking what percentage are 2 std. deviations above the mean, and that is about 2.5%.Normalcdf(2,6) is 2.28%. The 6 is large enough for the top number. One may use 1e99, but 6 works as well.
b. This is asking what percentage are between 1 and 2 sd, and that his 16-2.5 or about 13.5%. The actual amount can be found using 2ndVARS2 normal cdf (1,2) and is 13.59%.
c. z(0.30) is infvnorm(0.30,0,1) or -0.5244
-0.5244=(x-120)/20
-10.49=x-120
x=109.51 mm Hg
(medical note: unless this is done intra-arterially, the error is at least a few mm Hg from different ears or from even the same person a few seconds later hearing it).